OpenStudy (anonymous):

x(x-1)(x+2)>0 solve the inequality and the answer needs to be in interval notation and i have no idea what im doing

OpenStudy (anonymous):

this sucker changes sign at the zeros of the factors, namely at \[-2,0,1\]

OpenStudy (anonymous):

break the real line up in to four parts \[(-\infty,-2), (-2,0), (0,1), (1,\infty)\]

OpenStudy (anonymous):

since this is a polynomial of degree 3 with postive leading coefficient, the graph will be negative, positive, negative, positive in that order

OpenStudy (anonymous):

ok you lost me there a little bit how do i fins those numbers?

OpenStudy (anonymous):

these numbers \(-2,0,1\) ?

OpenStudy (anonymous):

yes are they just the ones that are in the equation?

OpenStudy (anonymous):

solve \(x(x-1)(x+2)=0\) in your head

OpenStudy (anonymous):

so when you graph \(y=x(x-1)(x+2)\) it will cross the \(x\) axis at \(-2\) and \(0\) and \(1\) as it crosses the \(x\) axis it goes from being negative (below) to positive (above) or vice versa

OpenStudy (anonymous):

so the answer in integral form would be -2< x<0 or x> 1

OpenStudy (anonymous):

that is not integral form or even interval form, that is as in inequality but yes that is right

OpenStudy (anonymous):

interval form is \((-2,0)\cup ((1,\infty)\)

OpenStudy (anonymous):

ic what does the u stand for?

OpenStudy (anonymous):

union

OpenStudy (anonymous):

i have an extra parentheses there, that is a typo

OpenStudy (anonymous):

ok so in order to find the answer you must find the solutions on the graph correct?

OpenStudy (anonymous):

no you can do that stupid "test point" business but it is a waste of time if you know that \(x(x-1)(x+2)>0\) is a synonym for "positive" and that positive is a synonym for "above the \(x\) axis" and if you know what a third degree polynomial looks like then you can do it in your head

OpenStudy (anonymous):

otherwise you could test a point, say \(x=-3\) and see that it is negative then you know it will be \(-,+,-,+\) on the intervals